OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

how can you find the slope

OpenStudy (anonymous):

lets do this together

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how can we find the slope

OpenStudy (anonymous):

put it in slope form

OpenStudy (anonymous):

(2, 2) and (6, 3) how can you get the slope from here?

OpenStudy (anonymous):

grr i lost

OpenStudy (anonymous):

\[m\frac{ y2-y1 }{ x2-x1 }\] now?

OpenStudy (anonymous):

6-3 --- 2-2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(2, 2) (x1, y1) (6, 3) (x2, y2) What are the x-values? What are the y-values?

OpenStudy (anonymous):

jazy do you like to teach her this one

OpenStudy (anonymous):

x is 2,6 y is 2,3

OpenStudy (anonymous):

ok now plug in to the equation above

OpenStudy (anonymous):

m= 3-2 ----- 6-2

OpenStudy (anonymous):

ok, now what you get?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now how can you write the equation in slope form?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

y=1x+4

OpenStudy (anonymous):

noo hmm

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

y = mx+b you got m and you have values for x and y so solve for b

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

(2, 2) y = mx + b 2 = (1/4)*2 + b solve for b

OpenStudy (anonymous):

@jazy im going to eat, please help her if she need more help

OpenStudy (anonymous):

2 = (1/4)*2 + b

OpenStudy (anonymous):

b = 3/2 right

OpenStudy (anonymous):

so now what

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

4x + 4y = –12

OpenStudy (anonymous):

x + 4y = 12

OpenStudy (anonymous):

can u tell me if im correct or help me or something

OpenStudy (anonymous):

3/2 is correct so now just plug in what you have y = mx + b plug in m and b and you are good

OpenStudy (anonymous):

y = 1/4x + 3/2 what do you multiply 1/4 with to get 1?